OpenStudy (anonymous):

A box contains 6 nuts, 8 bolts, and 4 screws. If 3 objects are selected in succession randomly, what is the probability of selecting a nut, then a bolt, then a screw, if replacement occurs each time?

OpenStudy (anonymous):

overall probability is the product of the individual probabilities. take each draw separately. first you want a nut. that is 6/(6+8+4) then that nut is put back. you then want a bolt. that is 8/(6+8+4) then that bolt is put back. you finally want a screw. that is 4/(6+8+4) take each of those three probabilities and multiply them together and you'll have your answer.

OpenStudy (anonymous):

i dont understand could you answer it for me. Math is my biggest weakness and i need this to graduate

OpenStudy (anonymous):

total objects=6+8+4=18 reqd. probability=6/18 *8/18 *4/18 solve and you get the answer.

OpenStudy (anonymous):

Sorry, i won't give the answers away. i am willing to help you figure it out, but i won't do the work for you.

OpenStudy (anonymous):

thanks alot yall